# Rate of epidemic spreading on complex networks

### v3

@@ Abstract
Equation (7) extends Eq. (1) to complex networks. Assuming exponential growth, $\dot{I}_{k^{\prime}}(u)\propto v_{k^{\prime}}e^{\Lambda t}$, we obtain for large time
@@
We shall use Eq. (8) as an implicit relation for the unknown $\Lambda$, and call $\Lambda_{\mathrm{EL}}$ the solution of the equation. We can also explicitly express $\Lambda$ from Eq. (8), but only in an approximate way. In particular, using the first two moments of $\psi(\tau)$, we obtain
@@
where $\sigma^{2}=\langle\tau^{2}\rangle-\langle\tau\rangle^{2}$, see sup . An alternative is to combine Eq. (8) with the Jensen inequality, that states that $\langle e^{-\Lambda\tau}\rangle\geq e^{-\Lambda\langle\tau\rangle}$. This leads to the lower bound $\Lambda\geq(\ln R)/\langle\tau\rangle$. This bound is saturated when $\psi(\tau)$ is a Dirac delta function.
Synthetic networks. The network Euler-Lotka equation successfully predicts the rate of epidemic spreading on synthetic complex networks, Fig. 2a and Fig. 2b. Our battery of tests includes a range of models characterized by different degree distribution, assortativity, and clustering. For each model, we use three different generation time distributions, see Fig. 2c. This choice qualitatively affect the epidemics. In particular, for peaked distributions, the exponential growth appears modulated by oscillations (Fig. 2a), since early generations of the epidemic are nearly synchronised. To extrapolate the leading exponential behavior from these curves, we fit them to a generalised logistic function
@@
where $N$ is the size of the network and $I_{0}$, $\Lambda_{0}$, and $\nu$ are free parameters, see sup . Expanding Eq. (10) for small $t$ we obtain $\log I(t)\approx\log I_{0}+t\Lambda_{0}[1-(I_{0}/N)^{\nu}]$. This leads us to define the finite-size growth rate $\Lambda_{\rm FIT}=\Lambda_{0}(1-\eta)$ and the finite-size parameter $\eta=\left(I_{0}/N\right)^{\nu}\in[0,1]$. The parameter $\eta$ is also a measure of deviations from an exponential behavior, see sup .
Figure 2: Network Euler-Lotka equation predicts the epidemic spreading rate on synthetic networks. (a) and (b): Average trajectory $I(t)$ over 500 realizations (red), $\Lambda_{\text{FIT}}$ (blue dot-dashed) and $\Lambda_{\text{EL}}$ (black dashed; based on $R_{\text{pert}}$, Eq. (6)) for Erdős-Rényi ($N=10^{6}$, $\overline{k}=4$) network with Gamma-distributed infection times and Watts-Strogatz network with Weibull-distributed infection times. (c) Solution of Eq. (8) compared with simulations for various networks of sizes up to $10^{6}$, see Ref. sup for details.
@@
When computing $\Lambda_{\mathrm{EL}}$, we set $R=R_{\text{pert}}$ as defined in Eq. (6) as the reproduction number. As expected, the predictions are considerably worse if we instead use $R_{0}$, see sup . The difference between $\Lambda_{\mathrm{FIT}}$ and $\Lambda_{\mathrm{EL}}$ is within $6\%$ for all networks, see Fig. 2c. The only exception is the Barabási-Albert model with parameter $m=1$ (average error $15\%$). One explanation for this discrepancy is that the exponential regime for this model is not well defined for $N\rightarrow\infty$, since $\bar{k^{2}}$ diverges in this limit. Another explanation is that, for $m=1$, $R_{\mathrm{pert}}$ does not accurately match the leading eigenvalue of the reproduction matrix (see Table 1).
@@
Figure 3: Network Euler-Lotka equation predicts the epidemic spreading rate on real-world networks. We obtained $I(t)$ by averaging $500$ simulations using log-normal infection times for each network from Refs. Clauset et al. (2016); Leskovec and Krevl (2014); Kunegis (2013) with at least $1000$ nodes sup . (a) and (b): Average trajectory $I(t)$ (red), $\Lambda_{\text{FIT}}$ (blue dot-dashed) and $\Lambda_{\text{EL}}$ (black dashed) for powergrid and PIN networks. (c) Exponential spreading rate $\Lambda_{\text{EL}}$ (dashed line) versus $\Lambda_{\text{FIT}}$ (markers). (d) Degree distribution tail exponent ($\gamma)$ and finite-size parameter ($\eta$) affect the estimation error $\varepsilon=2|\Lambda_{\text{FIT}}-\Lambda_{\rm EL}|/(\Lambda_{\text{FIT}}+\Lambda_{\rm EL})$. Each dot represents a network, dashed lines mark the scale-free region $2<\gamma\leq 3$. (e) and (f): Average degree $\bar{k}_{n}$ of infected individuals at $n$-th generation for (a) Erdős-Rényi ($N=10^{6}$, $\overline{k}=4$) and (b) Youtube (scale-free, $\gamma\approx 2.2$). Dashed lines represent power-law fits for $n\in[1,6]$.
@@
Real-world networks. We apply the network Euler-Lotka equation to a large set of social, biological, technological, and transportation networks. These real-world networks often exhibit more complex structures than synthetic networks, with tightly linked communities, strong degree-correlation and broad tails in the degree distribution. Although they constitute a challenging test, the network Euler-Lotka equation holds well, see Fig. 3a and Fig. 3b for examples.
More generally, we find that $\Lambda_{\text{EL}}$ well matches $\Lambda_{\text{FIT}}$ whenever the finite-size parameter $\eta$ detects a clear exponential regime, see Fig. 3c. Although some of these networks are highly clustered, using $R_{\mathrm{pert}}$ as expressed by Eq. (6) does not perform worse than $R$ sup . In Fig. 3c, we have omitted three road networks, for which the fit of Eq. (10) fails to converge. However, a function of the form $\dot{I}(t)\sim t^{\beta}$ well fits the epidemic trajectories in these cases sup , likely because these networks are embedded in a two-dimensional physical space (see Moore et al. (2024)).
@@
For networks in which $\eta<0.5$, the predictions of Eq. (8) using $R_{\text{pert}}$ are more accurate than those using $R_{0}$. However, for larger values of $\eta$, $R_{0}$ performs slightly better on average sup . Even in these cases, $R_{\mathrm{pert}}$ provides a better approximation of $R$ sup . The advantage of $R_{0}$ for large $\eta$ might be due to a compensatory effect: using $R_{0}$ tends to overestimate $R$, while the network Euler-Lotka equation tends to underestimate $\Lambda$ compared to $\Lambda_{\mathrm{FIT}}$.
@@
All of the networks where the prediction error is large are scale-free and characterized by a large value of $\eta$, see Fig. 3d. Here, scale-free means that the tail of the degree distribution scales as $k^{-\gamma}$, with $\gamma\leq 3$, see Voitalov et al. (2019) and sup . In contrast, the error is weakly affected by clustering sup . The error for scale-free networks can be explained by the early infection of hubs, which depletes the tail of the degree distribution. In networks that are not scale-free, we find that the average degree of infected individual is constant among generations as expected (Fig. 3e). In contrast, we find this average degree to decay as a power law in scale-free networks (Fig. 3f), leading to a non-steady spreading.
Discussion. In this Letter, we have related the exponential spreading rate of an epidemic on a complex network with the infectiousness and the network structure. Our theory accurately predicts the spreading rate on a wide range of synthetic and real-world networks, aside from scale-free networks where the exponential regime is hindered by finite-size effects.

### v4

@@ V Network Euler-Lotka equation
Equation (17) descends from the definitions of $M_{kk^{\prime}}$ and $\psi(\tau)$, and extends Eq. (1) to complex networks. Assuming exponential growth, $\dot{I}_{k^{\prime}}(u)\propto v_{k^{\prime}}e^{\Lambda t}$, we obtain for large time
@@
We shall use Eq. (18) as an implicit relation for the unknown $\Lambda$, and call $\Lambda_{\mathrm{EL}}$ the solution of the equation. In some cases, the equation can be explicitly solved. For example, in the Markovian limit (constant infectiousness $\lambda(t)=\lambda$ and $\psi(t)=\lambda e^{-\lambda t}$), Eq. (18) predicts that $\Lambda=\lambda(R-1)$. In general, we can only explicitly express $\Lambda$ from Eq. (18) in an approximate way. In particular, using the first two moments of $\psi(\tau)$, we obtain
@@
where $\sigma^{2}=\langle\tau^{2}\rangle-\langle\tau\rangle^{2}$, see Appendix B. An alternative is to combine Eq. (18) with the Jensen inequality, that states that $\langle e^{-\Lambda\tau}\rangle\geq e^{-\Lambda\langle\tau\rangle}$. This leads to the lower bound
$\Lambda\geq(\ln R)/\langle\tau\rangle\,.$ (20)
This bound is saturated when $\psi(\tau)$ is a Dirac delta function.
V.1 Synthetic networks
Figure 3: Network Euler-Lotka equation predicts the epidemic spreading rate on synthetic networks. (a) and (b): Average trajectory $I(t)$ over 500 realizations (red), $\Lambda_{\text{FIT}}$ (blue dot-dashed) and $\Lambda_{\text{EL}}$ (black dashed; based on $R_{\text{pert}}$, Eq. (6)) for Erdős-Rényi ($N=10^{6}$, $\overline{k}=4$) network with Gamma-distributed infection times and Watts-Strogatz network with Weibull-distributed infection times. (c) Solution of Eq. (18) compared with simulations for various networks of sizes up to $10^{6}$, see Appendix C and Cure et al. (2024).
The network Euler-Lotka equation predicts the rate of epidemic spreading on synthetic complex networks with remarkable accuracy, Figs. 3(a) and  3(b). Our battery of tests includes a range of models characterized by different degree distribution, assortativity, and clustering. For each model, we use three different generation time distributions $\psi(\tau)$, see Fig. 3(c). All three distributions are characterized by $p_{\psi}=1$; we shall consider the case $p<\psi<1$ in Sec. V.3. The choice of $\psi(\tau)$ qualitatively affects the epidemic spreading. In particular, for peaked distributions, the exponential growth appears modulated by oscillations [Fig. 3(a)], since early generations of the epidemic are nearly synchronized. To extrapolate the leading exponential behavior from these curves, we fit them into a generalized logistic function,
@@ V.1 Synthetic networks
which appears to well capture finite-size effects. In Eq. (21), $N$ the size of the network and $I_{0}$, $\Lambda_{0}$, and $\nu$ are free parameters. We fit the logarithm of $I(t)$ versus $t$ and discard the initial times at which $I(t)<\overline{k}$. The reason is that the first $\overline{k}$ infections should belong to the first generation, in which the average number of secondary infections per individual is $Z_{1}=\overline{k}$ rather than $R$. Although each simulation starts with a single infected individual, we consider $I_{0}$ as a free parameter to better capture the behavior of the epidemic at the early stages. Expanding Eq. (21) for small $t$, we obtain
$\log I(t)\approx\log I_{0}+t\Lambda_{0}[1-(I_{0}/N)^{\nu}]\,.$ (22)
This leads us to define the finite-size growth rate
$\Lambda_{\rm FIT}=\Lambda_{0}(1-\eta)\,,$ (23)
and the finite-size parameter
$\eta=\left(I_{0}/N\right)^{\nu}\in[0,1]\,,$ (24)
which quantifies the impact of finite-size effects on the exponential spreading rate.
@@
When computing $\Lambda_{\mathrm{EL}}$, we set $R=R_{\text{pert}}$ as defined in Eq. (6) as the reproduction number. As expected, the predictions are considerably worse if we instead use $R_{0}$, see Tables III-VIII in Supplementary Material sup and Appendix E. The difference between $\Lambda_{\mathrm{FIT}}$ and $\Lambda_{\mathrm{EL}}$ is within $6\%$ for all networks, see Fig. 3(c). The only exception is the Barabási-Albert model with parameter $m=1$ (average error $15\%$). One explanation for this discrepancy is that the exponential regime for this model is not well defined for $N\rightarrow\infty$, since $\bar{k^{2}}$ diverges in this limit. Another explanation is that, for $m=1$, $R_{\mathrm{pert}}$ does not accurately match the leading eigenvalue of the reproduction matrix (see Table 1).
@@
Figure 4: Network Euler-Lotka equation predicts the epidemic spreading rate on real-world networks. We obtained $I(t)$ by averaging $500$ simulations using log-normal infection times for each network from Refs. Clauset et al. (2016); Leskovec and Krevl (2014); Kunegis (2013) with at least $1000$ nodes sup . In all cases, we have set $p_{\psi}=1$. (a) and (b): Average trajectory $I(t)$ (red), $\Lambda_{\text{FIT}}$ (blue dot-dashed), and $\Lambda_{\text{EL}}$ (black dashed) for power grid and PIN networks. (c) Exponential spreading rate $\Lambda_{\text{EL}}$ (dashed line) versus $\Lambda_{\text{FIT}}$ (markers). (d) Degree distribution tail exponent ($\gamma)$ and finite-size parameter ($\eta$) affect the estimation error $\varepsilon=2|\Lambda_{\text{FIT}}-\Lambda_{\rm EL}|/(\Lambda_{\text{FIT}}+\Lambda_{\rm EL})$. Each dot represents a network, dashed lines mark the scale-free region $2<\gamma\leq 3$. (e) and (f) Average degree $\bar{k}_{n}$ of infected individuals at $n$-th generation for (a) Erdős-Rényi ($N=10^{6}$, $\overline{k}=4$) and (b) YouTube (scale-free, $\gamma\approx 2.2$). Dashed lines represent power-law fits for $n\in[1,6]$.
V.2 Real-world networks
@@
We apply the network Euler-Lotka equation to a large set of social, biological, technological, and transportation networks. These real-world networks often exhibit more complex structures than synthetic networks, with tightly linked communities, strong degree correlation, and broad tails in the degree distribution. Although they constitute a challenging test, the network Euler-Lotka equation holds well, see Figs. 4(a)  4(b) for examples. More generally, we find that $\Lambda_{\text{EL}}$ well matches $\Lambda_{\text{FIT}}$ whenever the finite-size parameter $\eta$ is small, see Fig. 4(c).
The fact that $\eta$ is small implies the presence of a clear exponential regime. To show that, we express the variation of the instantaneous exponential growth rate over a given timescale $\delta t$ as $d^{2}\ln I(t)/dt^{2}|_{t\rightarrow 0+}\delta t$. Substituting Eq. (21), this implies that the relative error on the exponent can be estimated as
$-\frac{\delta t}{\Lambda}\left.\frac{d^{2}\ln I(t)}{dt^{2}}\right|_{t\rightarrow 0+}=\nu\eta\Lambda_{0}\delta t\,,$ (25)
which is proportional to $\eta$. For example, in the case of the power grid network, we have $\eta=0.5$, $\nu=0.09$, and $\Lambda_{0}=0.35$. In this case, the relative error becomes of order one over a timescale $\delta t\sim(\nu\eta\Lambda_{0})^{-1}\approx 60$ days, which is before the epidemic has reached saturation, see Fig. 4(a). Thus the parameter $\eta$ also serves as a measure of deviations from an exponential behavior.
@@
Although some of these networks are highly clustered, using $R_{\mathrm{pert}}$ as expressed by Eq. (6) does not perform worse than $R$, see Tables in sup . In Fig. 4c, we have omitted three road networks, for which the fit of Eq. (21) fails to converge. However, a function of the form $\dot{I}(t)\sim t^{\beta}$ well fits the epidemic trajectories in these cases, see Appendix D. This is likely because these networks are embedded in a two-dimensional physical space, see Moore et al. (2024). For networks in which $\eta<0.5$, the predictions of Eq. (18) using $R_{\text{pert}}$ are more accurate than those using $R_{0}$. However, for larger values of $\eta$, $R_{0}$ performs slightly better on average. Even in these cases, $R_{\mathrm{pert}}$ provides a better approximation of $R$, see Appendix E. The advantage of $R_{0}$ for large $\eta$ might be due to a compensatory effect: using $R_{0}$ tends to overestimate $R$, while the network Euler-Lotka equation tends to underestimate $\Lambda$ compared to $\Lambda_{\mathrm{FIT}}$.
All of the networks where the prediction error is large are scale-free and characterized by a large value of $\eta$, see Fig. 4(d). Here, scale-free means that the tail of the degree distribution scales as $k^{-\gamma}$, with $\gamma\leq 3$. Specifically, a degree distribution is considered heavy-tailed or scale-free if it can be written in the form $P(k)=\ell(k)k^{-\gamma}$, where $\ell(k)$ is a slowly varying function satisfying $\lim_{k\rightarrow\infty}\ell(xk)/\ell(k)=1$, see Voitalov et al. (2019); Broido and Clauset (2019). This generalizes beyond pure power-law distributions, accounting for real-world data often better fit by distributions like lognormal or those associated with models such as the Barabási-Albert , which do not follow a pure power lawPósfai and Barabási (2016). We estimate the exponent $\gamma$ following the approach from Ref. Voitalov et al. (2019). The method estimates an extreme value index $\xi=(\gamma-1)^{-1}$ using kernel estimators Groeneboom et al. (2003); Voitalov (2018). If the kernel method fails, we resort to the method of moments. The specific values of $\gamma$ estimated for each real network using different methods are provided in the Supplementary Material sup .
@@
The error for scale-free networks can be explained by the early infection of hubs, which depletes the tail of the degree distribution. In networks that are not scale-free, we find that the average degree of infected individuals is constant among generations as expected [Fig. 4(e)]. In contrast, we find this average degree to decay as a power law in scale-free networks [Fig. 4(f)], leading to a nonsteady spreading. We also tested whether clustering affects the accuracy of the predicted spreading rate and found no significant effect, see Fig. 5.
Figure 5: Effect of degree distribution tail exponent ($\gamma)$, finite-size parameter ($\eta$) and clustering coefficient ($\alpha)$ on the estimation error $\varepsilon=2|\Lambda_{\rm FIT}-\Lambda_{\rm EL}|/(\Lambda_{\rm FIT}+\Lambda_{\rm EL})$. Each dot represents a network, dashed lines mark the scale-free region $2<\gamma\leq 3$. (a) Effect of $\gamma$ and $\eta$ on $\epsilon$ [identical to Fig. 4(d)]. (b) Effect of $\gamma$ and $\alpha$ on $\epsilon$.
V.3 Case $p_{\psi}<1$
We numerically test the case $p_{\psi}<1$ on a Watts–Strogatz network, which is clustered and has an almost vanishing correlation in the connected degrees ($r\approx 0)$. Our simulations are in excellent agreement with the predictions of Eq. (12), see Fig. 6, provided that $p_{\psi}$ is large enough so that the decimated network is well above its percolation threshold and therefore guaranteed to have a giant component.
Figure 6: Effect of non-normalized generation-time on the spreading rate. We generate a Watts–Strogatz network with parameter $k=6$ and rewiring probability $p=0.3$. We then simulate the average epidemic trajectory over $100$ trajectories using a Gamma generation-time distribution, imposing that the transmission via each link occurs with probability $p_{\psi}$. We repeat this for different values of $p_{\psi}$ and measure the empirical spreading rate (red markers). The black dashed line is the prediction of Eq. (12). The blue dashed line marks the spreading rate for $p_{\psi}=1$.
To test Eq. (IV.2), we simulate epidemics with a Gamma generation-time distribution while varying $p_{\psi}$ on two different networks. First, we consider a synthetic network that is disassortative and exhibits a heavy-tail in its degree distribution, see Fig. 7(a). Second, we consider the Amazon network, where nodes represent products and edges link commonly copurchased products Leskovec and Krevl (2014); Leskovec et al. (2007). The Amazon network presents assortativity, a heavy tail, and clustering. We find that our theory is in excellent agreement with the simulations provided that $p_{\psi}$ is not too small, as expected, see Fig. 7(b).
(a) Lognormal network
(b) Amazon network
Figure 7: Impact of the transmission probability on the epidemic spreading rate. For each value of $p_{\psi}$ (red markers), we simulate the average epidemic trajectory over $100$ trajectories with infection times that are Gamma distributed. The chosen networks are :(a) a configuration model in which the degree is lognormally distributed, (b) the Amazon network. The black dashed line represent the solution of the Euler-Lotka equation with $R=R^{\psi}_{\mathrm{pert}}$ given by Eq. (IV.2).
VI Weighted networks
To realistically model epidemic spreading, it is often necessary to assume that interactions between individuals are not always of the same type. For instance, infected individuals might spread the disease more rapidly to contacts in closed environments, such as workplaces, compared to outdoor spaces. To account for this, we generalize the network Euler-Lotka equation to networks in which edges are assigned a type $i$ with corresponding weight $w_{i}$ and transmission time distribution
[…]
In this case, the renewal equation becomes
[…]
where we have defined
$M_{kk^{\prime}}^{(i)}=Q(i|k,k^{\prime})M_{kk^{\prime}}$ (28)
where $Q(i|k,k^{\prime})$ is the fraction of edges with weight $w_{i}$ among the edges connecting nodes of degree $k$ and $k^{\prime}$. Since $\sum_{i}Q(i|k,k^{\prime})=1$ by construction, we have that $M_{kk^{\prime}}=\sum_{i}M_{kk^{\prime}}^{(i)}$. Assuming exponential growth, $\dot{I}_{k^{\prime}}(u)\propto v_{k^{\prime}}e^{\Lambda t}$ for large times, we obtain the weighted Euler-Lotka equation,
$v_{k}=\sum_{k^{\prime}}\left(\sum_{i}\big{\langle}e^{-\Lambda\tau}\big{\rangle}_{i}M_{kk^{\prime}}^{(i)}\right)v_{k^{\prime}},$ (29)
[…]
$\tilde{M}_{kk^{\prime}}(\Lambda)=\sum_{i}\big{\langle}e^{-\Lambda\tau}\big{\rangle}_{i}M_{kk^{\prime}}^{(i)}\,,$ (30)
the weighted Euler-Lotka equation imposes that the exponential growth rate is the value of $\Lambda$ such that the leading eigenvalue of $\tilde{M}_{kk^{\prime}}(\Lambda)$ is equal to one. The derivation can be readily extended to a continuous set of weights by replacing the sums over $i$ with integrals; in this case, the function $Q$ becomes a probability density over the weights.
Equation (29) successfully predicts the exponential growth rate for weighted networks constructed as combinations of different network models, see Fig. 8. In these examples, each network is constructed so that edges of each type are generated by a different network model.
Figure 8: Exponential spreading rate on weighted networks with two types of edges. Each network has $10^{5}$ nodes and two types of edges with weights $w_{1}$ and $w_{2}$. We run a simulation in which the generation time distribution for $w_{i}=1$ is Gamma distributed with mean $5$ and variance $3$. For different value of $w_{i}$, the generation time distribution is rescaled accordingly. (a) Edges with weight $w_{1}=1$ form a Watts–Strogatz network with parameter $k_{0}=6$ and $p=0.8$ and edges with weight $w_{2}=4$ form a Watts–Strogatz network with parameter $k_{0}=4$ and $p=0.4$. (b) Edges with weight $w_{1}=1$ form a Barabási-Albert network with parameter $m=3$ while edges with weight $w_{2}=12$ form a Watts–Strogatz network with parameter $k_{0}=6$ and $p=0.6$. (c) Edges with weight $w_{1}=2$ form a Erdős-Rényi network with mean degree $10$ while edges with weight $w_{2}=7$ form a Watts–Strogatz network with parameter $k_{0}=4$ and $p=0.3$. (d) Edges with weight $w_{1}=1.5$ form a Erdős-Rényi network with mean degree $10$ while edges with weight $w_{2}=4.1$ form a Barabási-Albert network with parameter $m=4$. The discrepancy between the solution of the weighted network Euler-Lotka equation (31) and the numerically computed spreading rate $\Lambda_{\rm FIT}$ is of $2\%$ in (a), $6\%$ in (b), less than $1\%$ in (c), and $2\%$ in (d).
We now consider the case in which each edge $e_{ij}$ has a random weight $w_{ij}$ which is identically and independently distributed for all edges. This case is particularly instructive, as it shows more clearly the formal relation between the weighted Euler-Lotka equation and the network Euler-Lotka equation introduced in Eq. (18).
We call $q(w)\equiv Q(i|k,k^{\prime})$ the probability of the weights, which in this case is independent of the degrees. Substituting into Eq. (29) we directly obtain
$\sum_{w}q(w)\left\langle e^{-\Lambda\tau}\right\rangle_{w}=\frac{1}{R},$ (31)
where $\langle\dots\rangle_{w}$ is an average over the distribution $\psi^{(w)}(\tau)$ for a given $w$ and $R$ is the leading eigenvalue of the matrix $M_{kk^{\prime}}$. Simulations for different network models and weight distributions support the predictions of Eq. (31), see Fig. 9.
(a) Barabási-Albert
(b) Erdős-Rényi
(c) Watts–Strogatz
Figure 9: Exponential spreading rate on networks with i.i.d weights. (a): Barabási-Albert network of size $10^{6}$ with parameter $m=7$. The infection times are lognormal with mean $5$ and variance $3$, the weights are $1$ with probability $p=0.9$ and $5$ with probability $p=0.1$. (b): Erdős-Rényi networks of size $10^{6}$ with mean degree $5$. the infection times are Weibull distributed with shape $2$ and scale $2$. The weights follow a power-law $1/w^{3}$ for $w\in[1,100]$. (c): Watts–Strogatz network of size $10^{6}$ with parameter $p=0.5$ and $k=4$. The infection times are Gamma distributed with mean $3$ and variance $1$. The weights are exponentially distributed with mean $3$. The discrepancy between the solution of the weighted network Euler-Lotka equation (Eq. (31) and the numerically computed spreading rate $\Lambda_{\rm FIT}$ is within $9\%$ in (a), less than $1\%$ in (b), and less than $1\%$ in (c).
VII Discussion